The sum of two integers is twenty-nine. Eight times the smaller integer is eleven more than five times the larger integer. Find the integers. large and small
Answer by richwmiller(17219)

a+b=29
8*a=11+5*b
a=12 b=17
check
96=11+5*17
96=11+85
ok